40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 3913     Entry time: Sat Nov 13 16:57:21 2010
Author: valera 
Type: Configuration 
Category: Electronics 
Subject: PRM Side OSEM transimpedance change 

Now that we have increased the range of the AA to +/- 10 V I have increased the PRM side OSEM transimpedance from 29 kV/A to 161 kV/A by changing the R64 in the satellite box. The first attached plot shows the ADC input spectrum before and after the change with analog whitening turned off. The PD voltage readback went up from 0.75 to 4.2 V. The second attached plot shows the sensor, ADC, and projected shot noise with analog whitening turned on and compensated digitally. The ADC calibration is 20 V/ 32768 cts. The PRM damping loops are currently disabled.

I checked for oscillation by looking at the monitor point at the whitening board. There was no obvious oscillation on a scope - the signal was 20 mV p-p on 1 us scale which was very similar to the LL channel.

Attachment 1: PRM-SD-ADC.pdf  92 kB  | Hide | Hide all
PRM-SD-ADC.pdf
Attachment 2: PRM-SD-Current.pdf  130 kB  | Hide | Hide all
PRM-SD-Current.pdf
ELOG V3.1.3-